# Hodge diamonds of the Landau–Ginzburg orbifolds

### 26 Feb 2024 v2

@@ 3.6. Symmetries and the Calabi–Yau condition
$\sum_{k=1}^{N}q_{k}=(1,\dots,1)E_{f}^{-1}\boldsymbol{1}=(1,\dots,1)(E_{f}^{T})^{-1}\boldsymbol{1}=\sum_{k=1}^{N}q_{k}^{T}.$ (25)
For a given $g$ assume $s$, such that $\overline{g}=E_{f}^{-1}s$ as in Proposition 23. None of $s_{k}=0$ because $N_{g}=0$. We have
$\mathrm{age}(g)$ $=\left((1,\dots,1)E_{f}^{-1}s\right)^{T}=s^{T}(E_{f}^{T})^{-1}\boldsymbol{1}=\sum_{k=1}^{N}s_{k}q_{k}^{T}.$ (26)
@@
$\sum_{k=1}^{N}s_{k}q_{k}^{T}\geq\sum_{k=1}^{N}q_{k}^{T}$
because every $s_{k}\geq 1$ and $q_{i}^{T}$ are all positive. Combining with Eq. (25) we get the inequality claimed. Moreover it’s obvious that the equality is only reached if $s_{k}=1$ for all $k$. This is equivalent to the fact that $g=j_{f}$.
Now let $f$ be star–shaped. We have $q_{1}^{T}=0$ and $q_{k}^{T}>0$ for $k=2,\dots,N$. By the same reasoning as above it’s enough to consider $\overline{g}=E_{f}^{-1}s$ with $s_{2}=\dots=s_{N}=1$. Then
@@
for some $s_{1}=1,\dots,a_{1}-1$. If $f$ defines an isolated singularity, it should have at least one summand $x_{2}^{r_{2}}\cdots x_{N}^{r_{N}}$ for some nonnegative $r_{2},\dots,r_{N}$. The quasihomogeneity and the $g$–invariance conditions on this summand give
$\sum_{k=2}^{N}\frac{r_{k}}{b_{k}}(1-\frac{1}{a_{1}})=1,\quad\sum_{k=2}^{N}\frac{r_{k}}{b_{k}}\frac{a_{1}-s_{1}}{a_{1}}\in{\mathbb{Z}}_{\geq 1}.$
@@
Remark 27.
The proposition above holds for any invertible polynomial without the Calabi–Yau condition too. However for noninvertible polynomial without Calabi–Yau condition the proposition does not hold in general. If particular for ${f=x_{1}^{10}+x_{1}(x_{2}^{2}+x_{3}^{2}+x_{4}^{2})}$ and $\overline{g}=(\frac{1}{5},\frac{2}{5},\frac{2}{5},\frac{2}{5})^{T}$ we have $\overline{q}=(\frac{1}{10},\frac{9}{20},\frac{9}{20},\frac{9}{20})$ and $\mathrm{age}(g)-\sum_{k=1}^{4}q_{k}=-\frac{1}{20}$.
4. The total space
@@
$\mathrm{Jac}(f):=\left.\raisebox{2.15277pt}{${\mathbb{C}}[x_{1},\dots,x_{N}]$}\mkern-3.0mu\middle/\mkern-3.0mu\raisebox{-2.15277pt}{$\left(\frac{\partial f}{\partial x_{1}},\dots,\frac{\partial f}{\partial x_{N}}\right)$}\right..$ (27)
It is a finite–dimensional ${\mathbb{C}}$–vector space whenever $f$ defines an isolated singularity. Call it Jacobian algebra of $f$ and set $\mu_{f}:=\dim_{\mathbb{C}}\mathrm{Jac}(f)$ — the Milnor number of $f$.
We will assume an additional convention:
for $\text{for the constant function}\ f=0\quad\text{set}\ \mathrm{Jac}(f):={\mathbb{C}},\ \mu_{f}:=1.$ (28)
4.1. Grading
The reduced weights $q_{1},\dots,q_{N}$ of $f$ define the ${\mathbb{Q}}$–grading on ${\mathbb{C}}[x_{1},\dots,x_{N}]$. Introduce the ${\mathbb{Q}}$–grading on $\mathrm{Jac}(f)$ by setting
$\deg([x_{1}^{\alpha_{1}}\cdots x_{N}^{\alpha_{N}}]):=\alpha_{1}q_{1}+\dots+\alpha_{N}q_{N}.$
Let $\phi_{1},\dots,\phi_{\mu}$ be the classes of monomials, generating $\mathrm{Jac}(f)$ as a ${\mathbb{C}}$–vector space. We say that $X\in\mathrm{Jac}(f)$ is of degree $\kappa$ if it’s expressed as a ${\mathbb{C}}$–linear combination of degree $\kappa$ elements $\phi_{\bullet}$.
@@ 4.1. Grading
$\mathrm{hess}(f):=\det\left(\frac{\partial^{2}f}{\partial x_{i}\partial x_{j}}\right)_{i,j=1,\dots,N}.$ (29)
@@
Proposition 28.
The maximal degree of a $\mathrm{Jac}(f)$–element is ${\widehat{c}}={\widehat{c}}(f):=\sum_{k=1}^{N}(1-2q_{k})$. Moreover we have
@@
See Section II of [AGV85]. ∎
4.2. Pairing
The algebra $\mathrm{Jac}(f)$ can be endowed with the ${\mathbb{C}}$–bilinear nondegenrate pairing $\eta_{f}$ called residue pairing (see [GH94, Chapter 5], [AGV85, Section 5.11]). The value $\eta_{f}([u],[v])$ is taken as the projection of $[u][v]$ to the top graded component $\mathrm{Jac}(f)_{\widehat{c}}$ divided by its generator $[\mathrm{hess}(f)]$. In particular, we have
$\eta_{f}([1],[\mathrm{hess}(f)])=1.$
Proposition 29.
@@ 4.2. Pairing
$\phi_{f,\beta}:\mathrm{Jac}(f)_{\beta}\cong(\mathrm{Jac}(f)_{{\widehat{c}}-\beta})^{\vee}$
$[p]\mapsto\eta_{f}([p],-),$
@@
See Section II of [AGV85]. ∎
4.3. The total space
For each $g\in{\mathrm{GL}}_{f}$ fix a generator of a one-dimensional vector space $\Lambda(g):=\bigwedge^{d_{g}}({\mathbb{C}}^{N}/\mathrm{Fix}(g))$. Denote it by $\xi_{g}$.
For $g\in G_{f}^{d}$ it is standard to choose the generator to be the wedge product of $x_{k}$ with $k\in I^{c}_{g}$ taken in increasing order.
Define ${\mathcal{B}}_{tot}(f)$ as the ${\mathbb{C}}$–vector spaces of dimension $\sum_{g\in{\mathrm{GL}}_{f}}\dim\mathrm{Jac}(f^{g})$
${\mathcal{B}}_{tot}(f):=\bigoplus_{g\in{\mathrm{GL}}_{f}}\mathrm{Jac}(f^{g})\xi_{g},$ (30)
Each direct summand $\mathrm{Jac}(f^{g})\xi_{g}$ will be called the $g$–th sector. We will write just ${\mathcal{B}}_{tot}$ when the polynomial is clear from the context.
Remark 30.
@@ 4.3. The total space
4.4. B-model group action
@@
$h\colon\mathrm{Fix}(g)\to\mathrm{Fix}(hgh^{-1}).$
and hence
$h\colon\Lambda(g)\to\Lambda(hgh^{-1}).$
Since we have fixed the generators $\xi_{\bullet}$, the latter map provides a constant $\rho_{h,g}\in{\mathbb{C}}^{*}$ such that
$h\left(\xi_{g}\right)=\rho_{h,g}\xi_{hgh^{-1}}.$
We have
$\rho_{h_{2},h_{1}gh_{1}^{-1}}\rho_{h_{1},g}=\rho_{h_{2}h_{1},g}.$ (31)
@@
$\rho_{h,g}=\prod_{\begin{subarray}{c}k=1,\dots,N\\ \lambda^{\prime}_{k}\neq 1\end{subarray}}\lambda_{k}$ (32)
We define the action of ${\mathrm{GL}}_{f}$ on ${\mathcal{B}}_{tot}$ by
$h^{*}([p({\bf x})]\xi_{g})=\rho_{h,g}[p(h^{-1}\cdot{\bf x})]\xi_{hgh^{-1}}.$ (33)
This is indeed a group action, i.e. $(h_{2}h_{1})^{*}=h_{2}^{*}\cdot h_{1}^{*}$. Indeed, using Eq. (31) we get
$(h_{2}h_{1})^{*}[p({\bf x})]\xi_{g}$ $=\rho_{h_{2}h_{1},g}[p((h_{2}h_{1})^{-1}\cdot{\bf x})]\xi_{(h_{2}h_{1})g(h_{2}h_{1})^{-1}}=$ (34)
$=\rho_{h_{2},h_{1}gh_{1}^{-1}}\rho_{h_{1},g}[p(h_{1}^{-1}h_{2}^{-1}\cdot{\bf x})]\xi_{h_{2}h_{1}gh_{1}^{-1}h_{2}^{-1}}=$ (35)
$=h_{2}^{*}\left(\rho_{h_{1},g}[p(h_{1}^{-1}\cdot{\bf x})]\xi_{h_{1}gh_{1}^{-1}}\right)=h_{2}^{*}h_{1}^{*}([p({\bf x})]\xi_{g}).$ (36)
@@
$h:\quad\xi_{g}\mapsto h^{*}(\xi_{g}):=\prod_{k\in I_{g}^{c}}h_{k}\,\cdot\xi_{g}$ (37)
Example 31.
Because $I_{\mathrm{id}}^{c}=I_{j_{f}}=\emptyset$ we have
$h^{*}(\xi_{\mathrm{id}})=\xi_{\mathrm{id}}\quad\text{and}\quad h^{*}(\xi_{j_{f}})=\det(h)\xi_{j_{f}}$
@@
${\mathcal{B}}_{tot,G}:=\bigoplus_{g\in G}\mathrm{Jac}(f^{g})\xi_{g}\subset{\mathcal{B}}_{tot}$
and define the B–model state space ${\mathcal{B}}(f,G)$ by
${\mathcal{B}}(f,G):=\left({\mathcal{B}}_{tot,G}\right)^{G}.$ (38)
Namely, the linear span of the ${\mathcal{B}}_{tot}$ vectors that are invariant with respect to the action of all elements of $G$.
Remark 32.
In the literature (see, for example, [M]) a different definition could be found where the sum is taken over the representatives of the conjugacy classes of $G$ and the invariants in each sector are taken with respect to the centralizer of the corresponding $g$. The two definitions are in fact equivalent in the same way as in [BI22, Proposition 42].
Example 33.
Let $f=x_{1}^{a_{1}}$ — the Fermat type polynomial. Assume $a_{1}=rm$ and consider $G$ to be generated by $g=(1/r)$. We have
${\mathcal{B}}_{tot}$ $={\mathbb{C}}\langle[1]\xi_{\mathrm{id}},[x_{1}]\xi_{\mathrm{id}},\dots,[x_{1}^{rm-2}]\xi_{\mathrm{id}}\rangle\oplus{\mathbb{C}}\langle[1]\xi_{g},\dots,[1]\xi_{g^{r-1}}\rangle.$
@@
$(g^{k})^{*}(\xi_{g^{l}})=\exp\left(2\pi i\cdot\frac{k}{r}\right)\xi_{g^{l}}.$
However $(g^{k})^{*}([x_{1}^{l}])=\exp\left(-2\pi i\cdot\frac{kl}{r}\right)[x_{1}^{l}]$ and the $G$–invariant monomials are $x_{1}^{rn}$ with $n\in{\mathbb{Z}}$. This gives
${\mathcal{B}}(f,G)$ $={\mathbb{C}}\langle[1]\xi_{\mathrm{id}},[x_{1}^{r}]\xi_{\mathrm{id}},\dots,[x_{1}^{r(m-1)}]\xi_{\mathrm{id}}\rangle.$
4.5. Bigrading
The following operators $q_{l},q_{r}:{\mathcal{B}}_{tot}\to{\mathbb{Q}}$ were first introduced in [IV90] giving the bigrading we use.

### 25 Mar 2024 v3

@@ 3.6 Symmetries and the Calabi–Yau condition
$\sum_{k=1}^{N}q_{k}=(1,\dots,1)E_{f}^{-1}\boldsymbol{1}=(1,\dots,1)\bigl{(}E_{f}^{\mathsf{T}}\bigr{)}^{-1}\boldsymbol{1}=\sum_{k=1}^{N}q_{k}^{\mathsf{T}}.$ (3.3)
For a given $g$ assume $s$, such that $\overline{g}=E_{f}^{-1}s$ as in Proposition 3.11. None of $s_{k}=0$ because $N_{g}=0$. We have
$\mathrm{age}(g)$ $=((1,\dots,1)E_{f}^{-1}s)^{\mathsf{T}}=s^{\mathsf{T}}\bigl{(}E_{f}^{\mathsf{T}}\bigr{)}^{-1}\boldsymbol{1}=\sum_{k=1}^{N}s_{k}q_{k}^{\mathsf{T}}.$
@@
$\sum_{k=1}^{N}s_{k}q_{k}^{\mathsf{T}}\geq\sum_{k=1}^{N}q_{k}^{\mathsf{T}}$
because every $s_{k}\geq 1$ and $q_{i}^{\mathsf{T}}$ are all positive. Combining with equation (3.3) we get the inequality claimed. Moreover it is obvious that the equality is only reached if $s_{k}=1$ for all $k$. This is equivalent to the fact that $g=j_{f}$.
Now let $f$ be star-shaped. We have $q_{1}^{\mathsf{T}}=0$ and $q_{k}^{\mathsf{T}}>0$ for $k=2,\dots,N$. By the same reasoning as above it is enough to consider $\overline{g}=E_{f}^{-1}s$ with $s_{2}=\dots=s_{N}=1$. Then
@@
for some $s_{1}=1,\dots,a_{1}-1$. If $f$ defines an isolated singularity, it should have at least one summand $x_{2}^{r_{2}}\cdots x_{N}^{r_{N}}$ for some nonnegative $r_{2},\dots,r_{N}$. The quasihomogeneity and the $g$-invariance conditions on this summand give
$\sum_{k=2}^{N}\frac{r_{k}}{b_{k}}\left(1-\frac{1}{a_{1}}\right)=1,\qquad\sum_{k=2}^{N}\frac{r_{k}}{b_{k}}\frac{a_{1}-s_{1}}{a_{1}}\in{\mathbb{Z}}_{\geq 1}.$
@@
Remark 3.15.
The proposition above holds for any invertible polynomial without the Calabi–Yau condition too. However for noninvertible polynomial without Calabi–Yau condition the proposition does not hold in general. If particular for ${f=x_{1}^{10}+x_{1}\bigl{(}x_{2}^{2}+x_{3}^{2}+x_{4}^{2}\bigr{)}}$ and $\overline{g}=\bigl{(}\frac{1}{5},\frac{2}{5},\frac{2}{5},\frac{2}{5}\bigr{)}^{\mathsf{T}}$ we have $\overline{q}=\bigl{(}\frac{1}{10},\frac{9}{20},\frac{9}{20},\frac{9}{20}\bigr{)}$ and $\mathrm{age}(g)-\sum_{k=1}^{4}q_{k}=-\frac{1}{20}$.
4 The total space
@@
$\mathrm{Jac}(f):=\left.\raisebox{2.15277pt}{${\mathbb{C}}[x_{1},\dots,x_{N}]$}\mkern-3.0mu\Big{/}\mkern-3.0mu\raisebox{-2.15277pt}{$\bigl{(}\frac{\partial f}{\partial x_{1}},\dots,\frac{\partial f}{\partial x_{N}}\bigr{)}$}\right..$
It is a finite-dimensional ${\mathbb{C}}$-vector space whenever $f$ defines an isolated singularity. Call it Jacobian algebra of $f$ and set $\mu_{f}:=\dim_{\mathbb{C}}\mathrm{Jac}(f)$ – the Milnor number of $f$.
We will assume an additional convention: for the constant function $f=0$ set $\mathrm{Jac}(f):={\mathbb{C}}$, $\mu_{f}:=1$.
4.1 Grading
The reduced weights $q_{1},\dots,q_{N}$ of $f$ define the ${\mathbb{Q}}$-grading on ${\mathbb{C}}[x_{1},\dots,x_{N}]$. Introduce the ${\mathbb{Q}}$-grading on $\mathrm{Jac}(f)$ by setting
$\deg\bigl{(}\bigl{[}x_{1}^{\alpha_{1}}\cdots x_{N}^{\alpha_{N}}\bigr{]}\bigr{)}:=\alpha_{1}q_{1}+\dots+\alpha_{N}q_{N}.$
Let $\phi_{1},\dots,\phi_{\mu}$ be the classes of monomials, generating $\mathrm{Jac}(f)$ as a ${\mathbb{C}}$-vector space. We say that $X\in\mathrm{Jac}(f)$ is of degree $\kappa$ if it is expressed as a ${\mathbb{C}}$-linear combination of degree $\kappa$ elements $\phi_{\bullet}$.
@@ 4.1 Grading
$\mathrm{hess}(f):=\det\left(\frac{\partial^{2}f}{\partial x_{i}\partial x_{j}}\right)_{i,j=1,\dots,N}.$
@@
Proposition 4.1.
The maximal degree of a $\mathrm{Jac}(f)$-element is ${\widehat{c}}={\widehat{c}}(f):=\sum_{k=1}^{N}(1-2q_{k})$. Moreover we have
@@
See [1, Section II]. ∎
4.2 Pairing
The algebra $\mathrm{Jac}(f)$ can be endowed with the ${\mathbb{C}}$-bilinear nondegenerate pairing $\eta_{f}$ called residue pairing (see [15, Chapter 5], [1, Section 5.11]). The value $\eta_{f}([u],[v])$ is taken as the projection of $[u][v]$ to the top graded component $\mathrm{Jac}(f)_{\widehat{c}}$ divided by its generator $[\mathrm{hess}(f)]$. In particular, we have $\eta_{f}([1],[\mathrm{hess}(f)])=1$.
Proposition 4.2.
@@ 4.2 Pairing
$\phi_{f,\beta}\colon\ \mathrm{Jac}(f)_{\beta}\cong(\mathrm{Jac}(f)_{{\widehat{c}}-\beta})^{\vee},\qquad[p]\mapsto\eta_{f}([p],-),$
@@
See [1, Section II]. ∎
4.3 The total space
For each $g\in{\mathrm{GL}}_{f}$, fix a generator of a one-dimensional vector space $\Lambda(g):=\bigwedge^{d_{g}}\bigl{(}{\mathbb{C}}^{N}/\mathrm{Fix}(g)\bigr{)}$. Denote it by $\xi_{g}$.
For $g\in G_{f}^{d}$, it is standard to choose the generator to be the wedge product of $x_{k}$ with $k\in I^{c}_{g}$ taken in increasing order.
Define ${\mathcal{B}}_{\rm tot}(f)$ as the ${\mathbb{C}}$-vector spaces of dimension $\sum_{g\in{\mathrm{GL}}_{f}}\dim\mathrm{Jac}(f^{g})$
${\mathcal{B}}_{\rm tot}(f):=\bigoplus_{g\in{\mathrm{GL}}_{f}}\mathrm{Jac}(f^{g})\xi_{g}.$ (4.1)
Each direct summand $\mathrm{Jac}(f^{g})\xi_{g}$ will be called the $g$-th sector. We will write just ${\mathcal{B}}_{\rm tot}$ when the polynomial is clear from the context.
Remark 4.3.
@@ 4.3 The total space
4.4 B-model group action
@@
$h\colon\ \mathrm{Fix}(g)\to\mathrm{Fix}\bigl{(}hgh^{-1}\bigr{)}\qquad\text{andhence}\qquad h\colon\ \Lambda(g)\to\Lambda\bigl{(}hgh^{-1}\bigr{)}.$
Since we have fixed the generators $\xi_{\bullet}$, the latter map provides a constant $\rho_{h,g}\in{\mathbb{C}}^{*}$ such that $h(\xi_{g})=\rho_{h,g}\xi_{hgh^{-1}}$. We have
$\rho_{h_{2},h_{1}gh_{1}^{-1}}\rho_{h_{1},g}=\rho_{h_{2}h_{1},g}.$ (4.2)
@@
$\rho_{h,g}=\prod_{\begin{subarray}{c}k=1,\dots,N\\ \lambda^{\prime}_{k}\neq 1\end{subarray}}\lambda_{k}.$
We define the action of ${\mathrm{GL}}_{f}$ on ${\mathcal{B}}_{\rm tot}$ by
$h^{*}([p({\bf x})]\xi_{g})=\rho_{h,g}\bigl{[}p\bigl{(}h^{-1}\cdot{\bf x}\bigl{)}\bigr{]}\xi_{hgh^{-1}}.$
This is indeed a group action, i.e., $(h_{2}h_{1})^{*}=h_{2}^{*}\cdot h_{1}^{*}$. Indeed, using equation (4.2) we get
$(h_{2}h_{1})^{*}[p({\bf x})]\xi_{g}$ $=\rho_{h_{2}h_{1},g}\bigl{[}p\bigl{(}(h_{2}h_{1})^{-1}\cdot{\bf x}\bigr{)}\bigr{]}\xi_{(h_{2}h_{1})g(h_{2}h_{1})^{-1}}$
$=\rho_{h_{2},h_{1}gh_{1}^{-1}}\rho_{h_{1},g}\bigl{[}p\bigl{(}h_{1}^{-1}h_{2}^{-1}\cdot{\bf x}\bigr{)}\bigr{]}\xi_{h_{2}h_{1}gh_{1}^{-1}h_{2}^{-1}}$
$=h_{2}^{*}\bigl{(}\rho_{h_{1},g}\bigl{[}p\bigl{(}h_{1}^{-1}\cdot{\bf x}\bigl{)}\bigl{]}\xi_{h_{1}gh_{1}^{-1}}\bigl{)}=h_{2}^{*}h_{1}^{*}([p({\bf x})]\xi_{g}).$
@@
$h\colon\ \xi_{g}\mapsto h^{*}(\xi_{g}):=\prod_{k\in I_{g}^{c}}h_{k}\,\cdot\xi_{g}.$
Example 4.4.
Because $I_{\mathrm{id}}^{c}=I_{j_{f}}=\varnothing$, we have
$h^{*}(\xi_{\mathrm{id}})=\xi_{\mathrm{id}}\qquad\text{and}\qquad h^{*}(\xi_{j_{f}})=\det(h)\xi_{j_{f}}$
@@
${\mathcal{B}}_{{\rm tot},G}:=\bigoplus_{g\in G}\mathrm{Jac}(f^{g})\xi_{g}\subset{\mathcal{B}}_{\rm tot}$
and define the B-model state space ${\mathcal{B}}(f,G)$ by ${\mathcal{B}}(f,G):=({\mathcal{B}}_{{\rm tot},G})^{G}$. Namely, the linear span of the ${\mathcal{B}}_{\rm tot}$ vectors that are invariant with respect to the action of all elements of $G$.
Remark 4.5.
In the literature (see, for example, [25]) a different definition could be found where the sum is taken over the representatives of the conjugacy classes of $G$ and the invariants in each sector are taken with respect to the centralizer of the corresponding $g$. The two definitions are in fact equivalent in the same way as in [3, Proposition 42].
Example 4.6.
Let $f=x_{1}^{a_{1}}$ – the Fermat type polynomial. Assume $a_{1}=rm$ and consider $G$ to be generated by $g=(1/r)$. We have
${\mathcal{B}}_{\rm tot}$ $={\mathbb{C}}\bigl{\langle}[1]\xi_{\mathrm{id}},[x_{1}]\xi_{\mathrm{id}},\dots,\bigl{[}x_{1}^{rm-2}\bigr{]}\xi_{\mathrm{id}}\bigr{\rangle}\oplus{\mathbb{C}}\langle[1]\xi_{g},\dots,[1]\xi_{g^{r-1}}\rangle.$
@@
$\bigl{(}g^{k}\bigr{)}^{*}(\xi_{g^{l}})=\exp\left(2\pi{\rm i}\cdot\frac{k}{r}\right)\xi_{g^{l}}.$
However $\bigl{(}g^{k}\bigr{)}^{*}\bigl{(}\bigl{[}x_{1}^{l}\bigr{]}\bigr{)}=\exp\bigl{(}-2\pi{\rm i}\cdot\frac{kl}{r}\bigr{)}\bigl{[}x_{1}^{l}\bigr{]}$ and the $G$-invariant monomials are $x_{1}^{rn}$ with $n\in{\mathbb{Z}}$. This gives
${\mathcal{B}}(f,G)$ $={\mathbb{C}}\bigl{\langle}[1]\xi_{\mathrm{id}},[x_{1}^{r}]\xi_{\mathrm{id}},\dots,\bigl{[}x_{1}^{r(m-1)}\bigr{]}\xi_{\mathrm{id}}\bigr{\rangle}.$
4.5 Bigrading
The following operators $q_{l},q_{r}\colon{\mathcal{B}}_{\rm tot}\to{\mathbb{Q}}$ were first introduced in [17] giving the bigrading we use.